OpenStudy (anonymous):

i have no idea how to do this GCF of term of polynomial 12x + 20

OpenStudy (shamil98):

gcf, 4(3x+5)

OpenStudy (anonymous):

correct

OpenStudy (shamil98):

The greatest common factor is a number that is the highest possible value that that binomial can be is factor of.

OpenStudy (shamil98):

same thing goes for trinomials and so on.

OpenStudy (phi):

they want you to find the same "thing" that is in both terms to find it, factor 12 into its prime factors 12= 2*6 = 2*2*3 factor 20 into 2*10= 2*2*5 so you have 2*2*3*x + 2*2*5 now what is in both terms? 2*2 ! factor out 2*2 (which is 4)